OpenStudy (anonymous):

3x+4y=4 and 3x+4=6 Are they parrell, perp, or neither?

OpenStudy (anonymous):

if it is \[3x+4y=4\] \[3x+4y=6\] they are parallel

OpenStudy (anonymous):

Linzee, First rewrite the equations in slope intercept form. y=mx+b you can do this by first subtracting the 3x on each side and then dividing by the number in front of the y variable on each side. If you follow the same procedure with the second equation you will see the slopes are the same and are therefore equal as was stated above.

OpenStudy (anonymous):

and if the slopes are equal, then they are parallel